OpenStudy (konradzuse):

Question about integral of arctan(2x)dx

OpenStudy (konradzuse):

So I got it down to the point where it's \[\int\limits \frac{2x}{1+4x^2}\]

OpenStudy (konradzuse):

or \[\int\limits \frac{2x}{1+(2x)^2}\]

OpenStudy (konradzuse):

now at first I thought I could cancel out the 2x, can I not do that?

OpenStudy (konradzuse):

because I got 1/2 ln|1+2x| and theirs was 1/4 ln|1+4x^2|

OpenStudy (lgbasallote):

let u = 2x du = 2dx \[\implies \frac 12 \int arctan (u) du\] let w = arctan u dw = 1/u^2 + 1 dv = du v = u \[wv - \int vdw\] \[\implies u \tan^{-1} u - \int \frac{u}{u^2 + 1}du\] now do trig sub  does that help?

OpenStudy (konradzuse):

you did it a much different way lol...

OpenStudy (konradzuse):

weird way boo iggy....

OpenStudy (konradzuse):

works tho....

OpenStudy (konradzuse):

Anyways dood...... Can I cancel the 2x from the (2x)^2 or not?

OpenStudy (lgbasallote):

\[\tan \theta = u\] \[\sec^2 \theta d\theta = du\] \[\sec\theta = \sqrt{u^2 + 1}\] \[\sec^2 \theta = u^2 + 1\] \[\implies \int \frac{\tan \theta \sec^2 \theta d\theta}{\sec^2 \theta}\] \[\implies \int \tan \theta d\theta\] \[\implies \ln | \sec \theta|\] \[\implies \ln | \sqrt{u^2 + 1}|\] okay this is sooo long

OpenStudy (lgbasallote):

no you cant

OpenStudy (konradzuse):

ok that's all I need to do.. and WTF ARE YOU DOING DOOD LOL.

OpenStudy (lgbasallote):

you'll get there....someday...lol

OpenStudy (konradzuse):

all I needed to know*

OpenStudy (lgbasallote):

and i typed all that =_=

OpenStudy (konradzuse):

leave me alone :'(.

OpenStudy (konradzuse):

Dude I said what I needed from the start :).

OpenStudy (lgbasallote):

really? must've missed it

OpenStudy (konradzuse):

so why can't I cancel it? 2x/(1+(2x)^2)?

OpenStudy (anonymous):

why not to try substitution let t=1+4x^2 dt=8xdx 1/8dt=xdx ??

OpenStudy (lgbasallote):

it's like \[\frac{a}{1+a^2}\] there's an addition...you can only cancel when it's multiplication like \[\frac{a}{ab}\]

OpenStudy (konradzuse):

Yup Sami that's what I was doing.

OpenStudy (lgbasallote):

anyway... how come you just got \[\int \frac{2x}{1 + (2x)^2} dx\]

OpenStudy (konradzuse):

But at first I tried to cancel, then I realized it's a du/u and then did it that way. When I went to check my answer they didn't cancel, and I didn't think about doing du/u before that.

OpenStudy (lgbasallote):

you used integration by parts right?

OpenStudy (konradzuse):

mhm

OpenStudy (lgbasallote):

you used u = arctan (2x) and dv = dx?

OpenStudy (konradzuse):

yessir.

OpenStudy (lgbasallote):

so where is the arctan in your solution?

OpenStudy (konradzuse):

the entire solution is xarctan(2x) - 1/4ln|1+4x^2|

OpenStudy (konradzuse):

the only part I was concerned with was why I couldn't cancel 2x from (2x)^2.. :'(?

OpenStudy (lgbasallote):

it was algebra

OpenStudy (lgbasallote):

algebra hindered you

OpenStudy (lgbasallote):

anyway...just so you know what we did are just same...that's why i told you you'll get there..in the end lol

OpenStudy (konradzuse):

IT aslways does bro I didn't have any teachers for all of middle school.

OpenStudy (konradzuse):

I know.... I'll just have to remember I cannot cancel like that... Idk why tho explain1

OpenStudy (lgbasallote):

can i use numbers to show you?